# MUGen: Multi-modal Music Understanding and Generation with the Power of Large Language Models

### v4

@@ 5.5 Subjective Evaluation for Music Generation
In order to provide a subjective assessment of our model’s music generation capabilities, we conduct a subjective evaluation involving 40 participants. A total of 20 questions are created for three subtasks: text-to-music (T2M), image-to-music (I2M), and video-to-music (V2M) generation. Each question has options generated by the models to be compared, which are randomly shuffled to avoid any preference bias from the testers. Since the sample size for subjective evaluation is relatively small, we use the interactive demo released by the authors of NExT-GPT to generate evaluation results for the T2M and I2M subtasks. However, for the V2M experiment, the NExT-GPT demo occasionally fail to function properly, so we decide to omit its comparison for V2M. The results are presented in Table 6. It shows that our proposed MUGen model consistently receive the highest preference among the testers for all three subtasks.
@@ 6 Conclusion and Future Work
This paper introduces the MUGen model, which utilizes a large language model (LLM) to achieve music understanding and multi-modal music generation within a unified framework. Furthermore, we present a comprehensive methodology for generating the datasets used to train our model. The experiments show that our proposed MUGen model outperforms or achieves SOTA performance in various tasks, including music understanding, music editing, and text/image/video-to-music generation. Our future work will focus on further enhancing the model’s fine-grained music understanding capabilities, as well as improving the correlation between generated music and input instructions.
@@ Appendix A Music Oriented Dataset Information
We generate 4 different datasets to train the MUGen model: MUCaps, MUImage, MUVideo and MUEdit datasets. The statistics of the datasets are given in Table 7. An example of each from the 4 datasets are shown in Figure 4.
@@ Appendix B Model Training
In this section, we detail the training strategy for the MUGen model along with parameters used for training.
@@ B.1 Model Training Strategy
The MUGen model adopts the adapter training strategy, implementing a three-step training regimen. In the first phase, all parameters, with the exception of those associated with the Multi-modal Understanding Adapters, undergo freezing. The training dataset is configured to incorporate the MUCaps dataset for music understanding, the COCO dataset for image comprehension, and the captions sourced from the MUVideo dataset for video understanding. During this training stage, the Cross Entropy Loss function is applied to compute the disparity between the caption generated by the LLaMA 2 model and the target caption corresponding to the input modality. This process is illustrated in Figure 5.
In the second training stage, the output projector is trained to generate conditional embeddings using input captions processed by the LLaMA 2 model. The LLaMA 2 model produces specialized audio tokens, denoted as [AUD] where (with as a hyperparameter representing the number of special audio tokens added to the LLaMA 2 model’s vocabulary) when processing input captions. The special audio tokens serve as signaling indicators, aiding the model in determining whether to generate text+music or solely text. In training, these audio tokens are added to the end of the text output in datasets requiring music output. During inference, if the MUGen model generates audio tokens, downstream music decoders (MusicGen/AudioLDM 2) will perform music generation, otherwise, solely text will be produced.
@@
In the final training stage, the LoRA training strategy is employed to train the LLaMA 2 model, concurrently fine-tuning the Multi-modal Understanding Adapter and Output Projection layer. This stage utilizes datasets including Alpaca, MusicQA, MUImage, MUVideo, and MUEdit. To signal the MUGen model to generate both music and text, the output text in MUImage, MUVideo, and MUEdit datasets is extended with the special audio tokens.
@@ Appendix C Model Evaluation
In this section, we elaborate on the datasets employed to assess the various capabilities of the MUGen model, followed by a discussion of the evaluation metrics utilized.
@@ C.1 Evaluation Datasets
For each of the training datasets generated—MUCaps, MUImage, MUVideo, and MUEdit—we create a corresponding evaluation set. The methodology employed for generating the evaluation dataset mirrors that of the training dataset generation. Detailed statistics for the evaluation datasets are provided in Table 8. However, for the MUEdit dataset, the evaluation set could not be utilized for model evaluation due to the unavailability of code bases and trained checkpoints for InstructME[27] and AUDIT[68]. Consequently, we resort to utilizing samples from InstructME’s demo website, which includes samples from both AUDIT and InstructME, to assess our model’s performance. For evaluating the MUGen’s music understanding capabilities we utilize the evaluation split of the MusicQA dataset.
@@ C.2 Evaluation Metrics
To assess music question answering, we adopt the metrics employed in [47], namely BLEU (B-U) [55], METEOR (M-R) [5], ROUGE (R-L) [43], and BERT-Score (BERT-S) [82]. These metrics are widely used for evaluating text generation. For all music generation tasks, we employ Fréchet Audio Distance (FAD) [37] and Kullback-Leibler divergence (KL), as these metrics are commonly utilized to assess the quality of generated audio. In addition to these general metrics, task-specific metrics are applied for each of the music generation tasks, namely Text-to-Music, Image-to-Music, Video-to-Music, and Music Editing.
@@
Using these evaluation metrics, we are able to evaluate the MUGen model against other state-of-the-art models for the different tasks.
@@ Appendix D Model Demonstration
In this section, we present screenshots of the MUGen model demo, illustrating various capabilities of the model.
Figures 8, 9, and 10 showcase the MUGen model’s ability to generate music directly from textual prompts and draw inspiration from images and videos, both with and without textual guidance. Figure 11 exemplifies MUGen’s proficiency in music editing guided by natural language prompts. Additionally, Figures 12, 13, and 14 illustrate the utilization of MUGen’s editing capabilities to further refine music generated from different modalities. Collectively, the MUGen model proves to be a robust framework for Music Understanding, Generation, and Editing.

### v5

@@ 5.5 Subjective Evaluation for Music Generation
In order to provide a subjective assessment of our model’s music generation capabilities, we conduct a subjective evaluation involving 40 participants. A total of 20 questions are created for three subtasks: text-to-music (T2M), image-to-music (I2M), and video-to-music (V2M) generation. Each question has options generated by the models to be compared, which are randomly shuffled to avoid any preference bias from the testers. Since the sample size for subjective evaluation is relatively small, we use the interactive demo released by the authors of NExT-GPT to generate evaluation results for the T2M and I2M subtasks. However, for the V2M experiment, the NExT-GPT demo occasionally fail to function properly, so we decide to omit its comparison for V2M. The results are presented in Table 6. It shows that our proposed M2UGen model consistently receive the highest preference among the testers for all three subtasks.
@@ 6 Conclusion and Future Work
This paper introduces the M2UGen model, which utilizes a large language model (LLM) to achieve music understanding and multi-modal music generation within a unified framework. Furthermore, we present a comprehensive methodology for generating the datasets used to train our model. The experiments show that our proposed M2UGen model outperforms or achieves SOTA performance in various tasks, including music understanding, music editing, and text/image/video-to-music generation. Our future work will focus on further enhancing the model’s fine-grained music understanding capabilities, as well as improving the correlation between generated music and input instructions.
@@ Appendix A Music Oriented Dataset Information
We generate 4 different datasets to train the M2UGen model: MUCaps, MUImage, MUVideo and MUEdit datasets. The statistics of the datasets are given in Table 7. An example of each from the 4 datasets are shown in Figure 4.
@@ Appendix B Model Training
In this section, we detail the training strategy for the M2UGen model along with parameters used for training.
@@ B.1 Model Training Strategy
The M2UGen model adopts the adapter training strategy, implementing a three-step training regimen. In the first phase, all parameters, with the exception of those associated with the Multi-modal Understanding Adapters, undergo freezing. The training dataset is configured to incorporate the MUCaps dataset for music understanding, the COCO dataset for image comprehension, and the captions sourced from the MUVideo dataset for video understanding. During this training stage, the Cross Entropy Loss function is applied to compute the disparity between the caption generated by the LLaMA 2 model and the target caption corresponding to the input modality. This process is illustrated in Figure 5.
In the second training stage, the output projector is trained to generate conditional embeddings using input captions processed by the LLaMA 2 model. The LLaMA 2 model produces specialized audio tokens, denoted as [AUD] where (with as a hyperparameter representing the number of special audio tokens added to the LLaMA 2 model’s vocabulary) when processing input captions. The special audio tokens serve as signaling indicators, aiding the model in determining whether to generate text+music or solely text. In training, these audio tokens are added to the end of the text output in datasets requiring music output. During inference, if the M2UGen model generates audio tokens, downstream music decoders (MusicGen/AudioLDM 2) will perform music generation, otherwise, solely text will be produced.
@@
In the final training stage, the LoRA training strategy is employed to train the LLaMA 2 model, concurrently fine-tuning the Multi-modal Understanding Adapter and Output Projection layer. This stage utilizes datasets including Alpaca, MusicQA, MUImage, MUVideo, and MUEdit. To signal the M2UGen model to generate both music and text, the output text in MUImage, MUVideo, and MUEdit datasets is extended with the special audio tokens.
@@ Appendix C Model Evaluation
In this section, we elaborate on the datasets employed to assess the various capabilities of the M2UGen model, followed by a discussion of the evaluation metrics utilized.
@@ C.1 Evaluation Datasets
For each of the training datasets generated—MUCaps, MUImage, MUVideo, and MUEdit—we create a corresponding evaluation set. The methodology employed for generating the evaluation dataset mirrors that of the training dataset generation. Detailed statistics for the evaluation datasets are provided in Table 8. However, for the MUEdit dataset, the evaluation set could not be utilized for model evaluation due to the unavailability of code bases and trained checkpoints for InstructME[27] and AUDIT[68]. Consequently, we resort to utilizing samples from InstructME’s demo website, which includes samples from both AUDIT and InstructME, to assess our model’s performance. For evaluating the M2UGen’s music understanding capabilities we utilize the evaluation split of the MusicQA dataset.
@@ C.2 Evaluation Metrics
To assess music question answering, we adopt the metrics employed in [47], namely BLEU (B-U) [55], METEOR (M-R) [5], ROUGEL (R-L) [43], and BERT-Score (BERT-S) [82]. These metrics are widely used for evaluating text generation. For all music generation tasks, we employ Fréchet Audio Distance (FAD) [37] and Kullback-Leibler divergence (KL), as these metrics are commonly utilized to assess the quality of generated audio. In addition to these general metrics, task-specific metrics are applied for each of the music generation tasks, namely Text-to-Music, Image-to-Music, Video-to-Music, and Music Editing.
@@
Using these evaluation metrics, we are able to evaluate the M2UGen model against other state-of-the-art models for the different tasks.
@@ Appendix D Model Demonstration
In this section, we present screenshots of the M2UGen model demo, illustrating various capabilities of the model.
Figures 8, 9, and 10 showcase the M2UGen model’s ability to generate music directly from textual prompts and draw inspiration from images and videos, both with and without textual guidance. Figure 11 exemplifies M2UGen’s proficiency in music editing guided by natural language prompts. Additionally, Figures 12, 13, and 14 illustrate the utilization of M2UGen’s editing capabilities to further refine music generated from different modalities. Collectively, the M2UGen model proves to be a robust framework for Music Understanding, Generation, and Editing.